OpenStudy (anonymous):

The population P of Texas (in thousands) from 1991 through 2000 can be modeled by: P=16,968e^0.019t where t=1 represents the year 1991. According to this model, when will the population reach 22 million?

jimthompson5910 (jim_thompson5910):

P is the population in thousands, so how do we represent 22 million?

OpenStudy (anonymous):

22,000?

jimthompson5910 (jim_thompson5910):

For example, if the population was 1000, then P = 1 if the population was 2000, then P = 2 etc etc

jimthompson5910 (jim_thompson5910):

good, multiply by 1000 to jump from thousands to millions

OpenStudy (anonymous):

so P=22,000?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that means you plug in P = 22000 and solve for t

jimthompson5910 (jim_thompson5910):

P=16,968e^(0.019t) 22000=16,968e^(0.019t) what's the next step?

OpenStudy (anonymous):

divide by 22000?

jimthompson5910 (jim_thompson5910):

that'll move over the 22000 to the right side

jimthompson5910 (jim_thompson5910):

we want to move over things that aren't t to the left side to isolate t

OpenStudy (anonymous):

I don't know how to do that haha

jimthompson5910 (jim_thompson5910):

16,968e^(0.019t) means 16,968 times e^(0.019t)

jimthompson5910 (jim_thompson5910):

so we first move over the 16,968 by undoing the multiplication

OpenStudy (anonymous):

so you would divide by 16, 968 then?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

giving you?

OpenStudy (anonymous):

1.296

jimthompson5910 (jim_thompson5910):

I'm getting 1.2965582272513, so that's roughly the same

jimthompson5910 (jim_thompson5910):

so that means e^(0.019t) = 1.2965582272513 what's next?

OpenStudy (anonymous):

Would you then figure out what e^0.019 is?

jimthompson5910 (jim_thompson5910):

what undoes exponentiation?

OpenStudy (anonymous):

take the log?

jimthompson5910 (jim_thompson5910):

correct, specifically the natural log (since that's base e)

jimthompson5910 (jim_thompson5910):

applying the natural log to both sides (ln...lowercase LN) gives e^(0.019t) = 1.2965582272513 ln(e^(0.019t)) = ln(1.2965582272513) 0.019t*ln(e) = ln(1.2965582272513) 0.019t*1 = ln(1.2965582272513) 0.019t = ln(1.2965582272513) hopefully you see which log rules I'm using above. If not, then let me know

jimthompson5910 (jim_thompson5910):

If you don't have any questions on that chunk of work, what's next?

OpenStudy (anonymous):

so then you would divide by 1.29etc to both sides?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you need to isolate t

jimthompson5910 (jim_thompson5910):

0.019t is the same as 0.019 times t

jimthompson5910 (jim_thompson5910):

you move that 0.019 over to isolate t by undoing whatever is being done to 0.019 and t

OpenStudy (anonymous):

sorry my internet connection is really bad right now. You would divide by .019 to isolate t by itself.

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

divide both sides by 0.019

jimthompson5910 (jim_thompson5910):

0.019t = ln(1.2965582272513) t = ln(1.2965582272513)/0.019 t = ???

OpenStudy (anonymous):

13.669117

jimthompson5910 (jim_thompson5910):

I'm getting the same

OpenStudy (anonymous):

so 13.7 years approximately. add that to 1991 and you'd get 2004.7?

jimthompson5910 (jim_thompson5910):

round that to the nearest whole year to get t = 14 so the population will happen in the middle of year 13, but it's definitely over 22 million on year 14

jimthompson5910 (jim_thompson5910):

I'd say 1991 + 14 = 2005

OpenStudy (anonymous):

So between 2004 and 2005

jimthompson5910 (jim_thompson5910):

yeah exactly

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

if they want a whole year, then go for 2005 since 2004 will fall short of 22 million

jimthompson5910 (jim_thompson5910):

yw

jimthompson5910 (jim_thompson5910):

oh wait

jimthompson5910 (jim_thompson5910):

t = 1 represents 1991

jimthompson5910 (jim_thompson5910):

so you add the result to 1990 not 1991

jimthompson5910 (jim_thompson5910):

1990 + 13 = 2003 1990 + 14 = 2004 so it's in between 2003 and 2004 (go with 2004 if they want a whole numbered year)